How much loan can a household afford if their gross annual income is RM80,000 with current mortgage rates at 6% p.a. compounded monthly over 20 years, monthly payments. (Hint: Use 25% of household income as the limit for payments.)
Answer #1

Max monthly payment = gross income/months in a year*limit of payment

=80000/12*0.25=1666.67

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
PV= 1666.67*((1-(1+ 6/1200)^(-20*12))/(6/1200))
PV = 232635.08 = Max loan amount